OpenStudy (anonymous):

Ruiz needs to buy a countertop for a kitchen. He calculated the area to be 16 square feet. The actual area is 15.4 square feet. What is Ruiz's percentage error?

OpenStudy (anonymous):

\[\frac{ 16-15.4 }{ 15.4 }\times 100=?\]

OpenStudy (anonymous):

3.9% ?